In 2 years from now you want to buy a new laptop. You expect the price of the laptop to be $1,750. How big are your equal weekly contributions to your savings account if you want to have enough money saved to buy a laptop in 2 years? Assume you make your first contribution to your savings account today and the last contribution at the end of 2 years, on the day when you buy the new laptop. Assume a year has 52 weeks. The interest rate is 12% (EAR).

Select one:

a. $18.6

b. $15.0

c. $18.8

d. $20.8

e. $13.4

f. $14.8

g. $16.8

h. $17.5

NOTE: Please show all the work, without using EXCEL (Step by step with equations) Thanks!
Answer #1

Answer: f. $ 14.8

Future value of an annuity = Annuity x Future Value Interest Factor of Annuity

Future value interest factor of an annuity = [ ( 1 + r ) n - 1 ] / r = [ ( 1.00231) 104 - 1 ] / 0.00231 = 117.4035

Since this is an annuity due, FVIF = 117.4035 x ( 1.00231) = 117.6747

Therefore, A x 117.6747 = $ 1,750

A ( weekly deposits) = $ 1,750 / 117.6747 = $ 14.87